A jokester tells you that he has a group of cows and chickens and that he counted 45 heads and 122 feet. How many cows and chickens does he have? (Do not use the strategy of making an equation or systems of equations for this item)


1
Expert's answer
2021-03-03T06:07:00-0500

Each cow has 1 head 4 feet and each chicken has 1 head and 2 feet.

Then the total number of animals is 45.

The number of cows is


"\dfrac{122-45(2)}{4-2}=16"

The number of chickens is


"45-16=29"